# Quantum concentration inequalities and equivalence of the thermodynamical ensembles: an optimal mass transport approach

### v1

@@ 1 Introduction
where each $\Lambda$ is a subset of the set of all the spins and each $h_{\Lambda}$ only acts on the spins in the corresponding $\Lambda$. The topic has originally been investigated for product states [2, 3, 4, 5] and later for high-temperature Gibbs states [6, 7] and states of spin lattices whose correlations decay exponentially with the distance [8]. Concentration inequalities for Gibbs states have been the key tool to prove the equivalence between the canonical and the microcanonical ensembles of quantum statistical mechanics and the weak version of the eigenstate thermalization hypothesis [9, 10, 11, 12, 7, 6].
@@
The best available concentration inequality for product states is [8, Theorem 5.3], and states that the concentration of local observables is exponential in the number of spins whenever each region $\Lambda$ in the sum (1.1) intersects only $O(1)$ other regions. Our Theorem 3.2 extends this result to the setting where each region $\Lambda$ contains $O(1)$ spins and the local norm of the observable is $O(1)$. We stress that our result covers also the case of observables with all-to-all interactions, e.g., where the sum in (1.1) runs over all the couples of spins, provided that the sum of the operator norms of the addends that act on any given spin is $O(1)$.
@@
The best available concentration inequalities for the states of a quantum spin lattice whose correlations decay exponentially with the distance are [8, Theorem 4.2] and [12, Eq. (S.17)], and state that the probability of deviations from the average decays as $\exp\left(-\Omega\left(n^{\frac{1}{2d+1}}\right)\right)$, where $n$ is the number of spins and $d>0$ depends on the geometry of the lattice, whenever the diameter of all the regions in the sum (1.1) is $O(1)$. Our Corollary 3.2 extends these results to the setting where each region $\Lambda$ contains $O(1)$ spins and the local norm of the observable is $O(1)$. Furthermore, while [8, Theorem 4.2] and [12, Eq. (S.17)] require a regular lattice, our results do not require any regular structure nor the notion of neighboring spins, but only require a distance on the set of the spins such that the size of each metric ball is $O(r^{d})$, where $r$ is the radius (Assumption 2.1).
@@
The second part of the paper is devoted to the equivalence between the ensembles of quantum statistical mechanics. It has long been known that a concentration inequality for an observable measured on a Gibbs state implies an upper bound in the difference between the expectation value of that observable on the Gibbs and on any microcanonical states with sufficiently close average energy [9, 10, 11, 12, 7, 6], thus bounding the distinguishability between the Gibbs and the microcanonical states and proving the equivalence between the canonical and microcanonical ensembles for the observables that exhibit concentration. Such link between concentration and equivalence of the ensembles has been connected to the quantum theory of optimal mass transport [13, 14, 15, 5]. In this context, a quantum generalization of the Lipschitz constant has been defined to quantify the maximum dependence of an observable on any given spin, and concentration inequalities have been proved for observables with $O(1)$ Lipschitz constant measured on weakly correlated states [6]. The distinguishability between quantum states with respect to observables with unit Lipschitz constant has been quantified by the quantum Wasserstein distance of order $1$, or quantum $W_{1}$ distance [5, 16]. Such concentration has been proved to be linked to transportation-cost inequalities that provide an upper bound to the $W_{1}$ distance between a generic state and a Gibbs state in terms of the square root of their relative entropy, and the equivalence between the canonical and microcanonical ensembles has been proved via such transportation-cost inequalities [6, Proposition 12]. It has further been proved that, whenever a Gibbs state satisfies such transportation-cost inequality, it is close with respect to the $W_{1}$ distance to any state that has approximately the same entropy, i.e., such that the difference between the entropy of the two states is $o(n)$, where $n$ is the number of spins.
@@ 2 Setup
$\left|\left\langle h_{\Lambda_{1}}\,h_{\Lambda_{2}}\right\rangle-\left\langle h_{\Lambda_{1}}\right\rangle\left\langle h_{\Lambda_{2}}\right\rangle\right|\leq C\,\exp\left(-\frac{\mathsf{d}(\Lambda_{1},\Lambda_{2})}{\xi}\right)\,,$ (2.5)
@@ 3.1 The local norm
The local norm (3.1) is similar to the norm considered in [7, Eq. (4)]. The novelty of (3.1) with respect to [7, Eq. (4)] is that, while [7, Eq. (4)] considers only a fixed decomposition of $H$ of the form (1.1), (3.1) optimizes over all the possible decompositions.
@@ 3.2 Product states
The function $F$ is strictly increasing with $F(0)=0$, then it is strictly positive for any $x>0$.
@@ 5 Equivalence of the ensembles of quantum statistical mechanics
In this paper, we consider the equivalence of the statistical mechanical ensembles and the weak ETH from the perspective of the $k$-local quantum $W_{1}$ distance. The closeness in such distance implies closeness of the expectation values of all the $k$-local observables with $O(1)$ local norm. We stress that such observables do not need to be geometrically local, since they can contain terms acting on spins at arbitrary distance. Our results are based on the transportation-cost inequality Theorem 4.1 and are similar in spirit to the results of [6, Sec. 8], which considers the equivalence of the ensembles with respect to the quantum $W_{1}$ distance of [5] employing transportation-cost inequalities for such distance. The key difference between the present paper and Ref. [6] is that the $k$-local quantum $W_{1}$ distance that we employ is weaker than the quantum $W_{1}$ distance of [6]. As a consequence, our Theorem 4.1 requires only exponentially decaying correlations. On the contrary, the transportation-cost inequalities employed in [6] require a Hamiltonian that is the sum of local commuting terms (the subsequent Ref. [29] has proved transportation-cost inequalities for the quantum $W_{1}$ distance under the hypothesis of exponential decay of the conditional mutual information, which is however still stronger than exponential decay of correlations). Therefore, we are able to prove the equivalence of the ensembles for a significantly larger class of Hamiltonians than the class for which the results of [6, Sec. 8] apply, at the price of a weaker notion of equivalence. We also stress that, contrarily to the results of Refs. [22, 12], we do not need to restrict to regular lattices, and our results do not require any regularity nor the notion of neighboring spins, but only the upper bound Assumption 2.1 to the growth of the size of the metric balls with the distance.
@@ 6 Conclusions
The main open problem we point out concerns the strong ETH, stating that the energy eigenstate themselves are close to the Gibbs states with the same average energy with respect to a suitable class of observables. From [6, Sec. 8], the strong ETH cannot be captured by the quantum $W_{1}$ distance of [5]. Indeed, the von Neumann entropy per spin is continuous with respect to the $W_{1}$ distance per spin [30], therefore no pure state can be close to any state with large entropy such a Gibbs state. On the contrary, we do not expect such entropy continuity to hold for the $k$-local quantum $W_{1}$ distance, and therefore there is hope that this distance can capture the strong ETH.

### v2

@@ Abstract
keywords:
Quantum spin systems, concentration inequalities, quantum Wasserstein distance, canonical and microcanonical ensembles
@@ 1 Introduction
where each $\Lambda$ is a subset of the set of all the spins and each $h_{\Lambda}$ only acts on the spins in the corresponding $\Lambda$. The topic has originally been investigated for product states [2, 3, 4, 5, 6] and later for high-temperature Gibbs states [7, 8], time-evolved product states [9], and states of spin lattices whose correlations decay exponentially with the distance [10]. Concentration inequalities for Gibbs states have been the key tool to prove the equivalence between the canonical and the microcanonical ensembles of quantum statistical mechanics and the weak version of the eigenstate thermalization hypothesis [11, 12, 13, 14, 8, 7].
@@
The best available concentration inequality for product states is [10, Theorem 5.3], and states that the concentration of local observables is exponential in the number of spins whenever each region $\Lambda$ in the sum (1.1) intersects only $O(1)$ other regions. Our Theorem 3.2 extends this result to the setting where each region $\Lambda$ contains $O(1)$ spins and the local norm of the observable is $O(1)$. We stress that our result covers also the case of observables with all-to-all interactions, e.g., where the sum in (1.1) runs over all the couples of spins, provided that the sum of the operator norms of the addends that act on any given spin is $O(1)$.
@@
The best available concentration inequalities for the states of a quantum spin lattice whose correlations decay exponentially with the distance are [10, Theorem 4.2] and [14, Eq. (S.17)], and state that the probability of deviations from the average decays as $\exp\left(-\Omega\left(n^{\frac{1}{2d+1}}\right)\right)$, where $n$ is the number of spins and $d>0$ depends on the geometry of the lattice, whenever the diameter of all the regions in the sum (1.1) is $O(1)$. Our Corollary 3.2 extends these results to the setting where each region $\Lambda$ contains $O(1)$ spins and the local norm of the observable is $O(1)$. Furthermore, while [10, Theorem 4.2] and [14, Eq. (S.17)] require a regular lattice, our results do not require any regular structure nor the notion of neighboring spins, but only require a distance on the set of the spins such that the size of each metric ball is $O(r^{d})$, where $r$ is the radius (Assumption 2.1).
@@
The second part of the paper is devoted to the equivalence between the ensembles of quantum statistical mechanics. It has long been known that a concentration inequality for an observable measured on a Gibbs state implies an upper bound in the difference between the expectation value of that observable on the Gibbs and on any microcanonical states with sufficiently close average energy [11, 12, 13, 14, 8, 7], thus bounding the distinguishability between the Gibbs and the microcanonical states and proving the equivalence between the canonical and microcanonical ensembles for the observables that exhibit concentration. Such link between concentration and equivalence of the ensembles has been connected to the quantum theory of optimal mass transport [15, 16, 17, 6]. In this context, a quantum generalization of the Lipschitz constant has been defined to quantify the maximum dependence of an observable on any given spin, and concentration inequalities have been proved for observables with $O(1)$ Lipschitz constant measured on weakly correlated states [7]. The distinguishability between quantum states with respect to observables with unit Lipschitz constant has been quantified by the quantum Wasserstein distance of order $1$, or quantum $W_{1}$ distance [6, 18]. Such concentration has been proved to be linked to transportation-cost inequalities that provide an upper bound to the $W_{1}$ distance between a generic state and a Gibbs state in terms of the square root of their relative entropy, and the equivalence between the canonical and microcanonical ensembles has been proved via such transportation-cost inequalities [7, Proposition 12]. It has further been proved that, whenever a Gibbs state satisfies such transportation-cost inequality, it is close with respect to the $W_{1}$ distance to any state that has approximately the same entropy, i.e., such that the difference between the entropy of the two states is $o(n)$, where $n$ is the number of spins.
@@ 2 Setup
$\left|\left\langle h_{\Lambda_{1}}\,h_{\Lambda_{2}}\right\rangle-\left\langle h_{\Lambda_{1}}\right\rangle\left\langle h_{\Lambda_{2}}\right\rangle\right|\leq C\left\|h_{\Lambda_{1}}\right\|\left\|h_{\Lambda_{2}}\right\|\exp\left(-\frac{\mathsf{d}(\Lambda_{1},\Lambda_{2})}{\xi}\right)\,,$ (2.5)
@@ 3.1 The local norm
The local norm (3.1) is similar to the norm considered in [8, Eq. (4)]. The novelty of (3.1) with respect to [8, Eq. (4)] is that, while [8, Eq. (4)] considers only a fixed decomposition of $H$ of the form (1.1), (3.1) optimizes over all the possible decompositions.
@@ 3.2 Product states
The function $F$ in (3.26) is strictly increasing with $F(0)=0$, then it is strictly positive for any $x>0$.
Remark 3.5.
The function $F$ in (3.26) is defined to keep the bound as tight as possible. By definition of $F$, the maximizer $s^{*}$ of $f(s)=s(x-e^{s}+1+s/2)$ is a function of $x\geq 0$ which does not admit an explicit form in terms of elementary functions. However, one can suppose to approximate $s^{*}$ by a suitable elementary function in order to explicitly obtain a function $\widetilde{F}:[0,+\infty)\rightarrow[0,+\infty)$, with $\widetilde{F}(x)\leq F(x)$ for every $x\geq 0$, which provides a suboptimal but more explicit bound in Equation 3.26. For any fixed $x\geq 0$, the optimum $s^{*}$ of $f$ satisfies:
$x=(e^{s^{*}}-1)(s^{*}+1).$
A rough approximation of the inverse of $x(s^{*})=(e^{s^{*}}-1)(s^{*}+1)$ is $\hat{s}^{*}(x)=\log(x+1)$ that provides the suboptimal version of $F$ by:
$\widetilde{F}(x)=f(\hat{s}^{*}(x))=\log(x+1)\left[x-e^{\log(x+1)}+1+\frac{1}{2}\log(x+1)\right]=\frac{1}{2}\log^{2}(x+1)$
Then, we get a worse but more explicit bound than the one given in Equation 3.26:
$\mathbb{P}\left(H\geq\left\langle H\right\rangle+na\right)\leq\exp\left[-\frac{n}{2k^{2}}\,\log^{2}\left(\frac{ak}{\|H\|_{\textnormal{loc}}}+1\right)\right],$ (3.29)
so we can easily note that the concentration weakens as the locality $k$ of $H$ increases.
3.2.1 Comparison with [6, Corollary 3]
Ref. [6] proves the following concentration inequality for the maximally mixed state:
Theorem 3.3 ([6, Corollary 3]).
Let $\omega$ be the maximally mixed state of $n$ spins. Then, for any $H\in\mathcal{O}_{[n]}$ and any $a>0$ we have
$\mathbb{P}\left(H\geq\left\langle H\right\rangle+na\right)\leq\exp\left[-\frac{2\,n\,a^{2}}{\left\|H\right\|_{L}^{2}}\right],$ (3.30)
where
$\left\|H\right\|_{L}=2\max_{x\in[n]}\min\left\{\left\|H-H_{[n]\setminus\{x\}}\otimes\mathbb{I}_{x}\right\|:H_{[n]\setminus\{x\}}\in\mathcal{O}_{H_{[n]\setminus\{x\}}}\right\}$ (3.31)
is the quantum Lipschitz constant of $H$.
For the maximally mixed state, [6, Corollary 3] is stronger than Theorem 3.2. Indeed, both results imply a concentration that is exponential in $n$ whenever $\left\|H\right\|_{L}$ and $\left\|H\right\|_{\mathrm{loc}}$, respectively, are $O(1)$. From [19, Proposition II.5] we have $\left\|H\right\|_{L}\leq\left\|H\right\|_{\mathrm{loc}}$, hence [6, Corollary 3] guarantees an exponential concentration for a larger family of Hamiltonians. However, the fundamental improvement of Theorem 3.2 is that it is valid not only for the maximally mixed state but for any product state, including pure product states. The validity of (3.30) for generic product states remains an open challenge.
@@ 3.3 Quantum states with finite correlation length
Remark 3.7.
We can write a concentration inequality that is expressed in terms of elementary functions and is more explicit than (3.47), at the cost of worsening the bound. In the proof of Corollary 3.2, the inequality (3.50) is obtained by optimizing over $s$. However, we can choose the approximate optimum $\hat{s}^{*}(x)=\log(x+1)$ as done in Remark 3.5. Leaving the selection of the value of $l$ given by (3.52) unchanged, the resulting bound is:
$\mathbb{P}(H\geq na)\leq(C+1)\exp\left[-\frac{n^{\frac{1}{2d+1}}\log^{2}\left(\frac{ak}{\|H\|_{\textnormal{loc}}}+1\right)}{2k^{2}A^{2}\left[\frac{\xi}{kA^{2}}\left(k\,\log\left(\frac{ak}{\|H\|_{\textnormal{loc}}}+1\right)+\frac{1}{2}\log^{2}\left(\frac{ak}{\|H\|_{\textnormal{loc}}}+1\right)\right)\right]^{\frac{2d}{2d+1}}}\right].$ (3.54)
3.3.1 Comparison with previous concentration bounds
Our concentration bound for states with exponential decay of correlations, Corollary 3.2, can be compared with existing results. In particular, under the same assumption on the correlation decay, in [14], the authors provide the following inequality:
$\mathbb{P}(H\geq\langle H\rangle+na)\geq\min\left(1,(e+3e\xi)\max\left(e^{-(\frac{na^{2}}{c})^{\frac{1}{d+1}}},e^{-\frac{na^{2}}{c^{\prime}k^{d}}}\right)\right),$ (3.55)
where $c$ and $c^{\prime}$ are constant depending on $d$ and $\xi$, and $k$ is the locality of $H$. A similar order of decay is provided by [10, Theorem 4.2]. However, these works have the requirement that the diameter of the regions, where the local terms of the observables act, is $O(1)$ and the spin lattice presents a regular structure.
In [7], there are some results about Gaussian concentration of the values of local observables. In particular [7, Theorem 7] provides a concentration bound for states satisfying a transportation cost inequality, defined by the quantum $W_{1}$ distance as defined in [6]. In the particular case where the state commutes with the considered observable $H$, the concentration assumes the form:
$\mathbb{P}(H\geq\langle H\rangle+na)\leq 2\exp\left(-\frac{(na)^{2}}{c\|H\|^{2}_{L}}\right)\,,$ (3.56)
where the constant $c$ is linked to a transportation cost inequality satisfied by the Gibbs state and is proved to be $O(n)$ at high enough temperature. However, [7, Theorem 7] requires that $H$ is a sum of commuting local terms.
In Table 1, we compare both the concentration bounds provided by Theorem 3.2 and Corollary 3.2 with some known results. In particular we highlight the assumptions under which the bounds are derived in order to stress the novelty of our concentration inequalities.
Concentration Quantum states Assumptions Ref.
$e^{-\Omega(n)}$ Product states $\|H\|_{\mathrm{loc}}$ is $O(1)$ Our result
Any $\Lambda$ contains $O(1)$ spins (Theorem 3.2)
$e^{-\Omega(n)}$ Maximally mixed state $\|H\|_{L}$ is $O(1)$ [18, Corollary 3]
$e^{-\Omega(n)}$ Product states Regions overlap is $O(1)$ [10, Theorem 5.2]
$\exp\left(-\Omega\left(n^{\frac{1}{2d+1}}\right)\right)$ States with finite $\|H\|_{\mathrm{loc}}$ is $O(1)$ Our result
correlation length Any $\Lambda$ contains $O(1)$ spins (Corollary 3.2)
$\exp\left(-\Omega\left(n^{\frac{1}{2d+1}}\right)\right)$ States with finite Diameter of $\Lambda$ is $O(1)$ [14, Equation S.17]
correlation length Regular spin lattice [10, Theorem 4.2]
$e^{-\Omega(n)}$ High-temperature states Commuting local terms in $H$ [7, Theorem 7]
Table 1: In the table, we report the concentration bounds presented in this paper and some previous known results. The first column reports the Bound over $\mathbb{P}(H\geq\langle H\rangle+na)$, the second one specifies the considered quantum states, and the third column contains the further assumptions under which the bounds were derived. The regions $\Lambda$ are those involved in the definition of the local observable $H$.
@@ 5 Equivalence of the ensembles of quantum statistical mechanics
In this paper, we consider the equivalence of the statistical mechanical ensembles and the weak ETH from the perspective of the $k$-local quantum $W_{1}$ distance. The closeness in such distance implies closeness of the expectation values of all the $k$-local observables with $O(1)$ local norm. We stress that such observables do not need to be geometrically local, since they can contain terms acting on spins at arbitrary distance. Our results are based on the transportation-cost inequality Theorem 4.1 and are similar in spirit to the results of [7, Sec. 8], which considers the equivalence of the ensembles with respect to the quantum $W_{1}$ distance of [6] employing transportation-cost inequalities for such distance. The key difference between the present paper and Ref. [7] is that the $k$-local quantum $W_{1}$ distance that we employ is weaker than the quantum $W_{1}$ distance of [7]. As a consequence, our Theorem 4.1 requires only exponentially decaying correlations. On the contrary, the transportation-cost inequalities employed in [7] require a Hamiltonian that is the sum of local commuting terms (the subsequent Ref. [30] has proved transportation-cost inequalities for the quantum $W_{1}$ distance under the hypothesis of exponential decay of the conditional mutual information, which is however still stronger than exponential decay of correlations). Therefore, we are able to prove the equivalence of the ensembles for a significantly larger class of Hamiltonians than the class for which the results of [7, Sec. 8] apply, at the price of a weaker notion of equivalence. We also stress that, contrarily to the results of Refs. [11, 14], we do not need to restrict to regular lattices, and our results do not require any regularity nor the notion of neighboring spins, but only the upper bound Assumption 2.1 to the growth of the size of the metric balls with the distance.
@@ 6 Conclusions
The main open problem we point out concerns the strong ETH, stating that the energy eigenstate themselves are close to the Gibbs states with the same average energy with respect to a suitable class of observables. From [7, Sec. 8], the strong ETH cannot be captured by the quantum $W_{1}$ distance of [6]. Indeed, the von Neumann entropy per spin is continuous with respect to the $W_{1}$ distance per spin [31], therefore no pure state can be close to any state with large entropy such a Gibbs state. On the contrary, we do not expect such entropy continuity to hold for the $k$-local quantum $W_{1}$ distance, and therefore there is hope that this distance can capture the strong ETH.
@@ Acknowledgements
Declarations
  • Conflict of interest: All authors have no conflicts of interest.
  • Data availability: The authors declare that the present manuscript has no associated data.